Currently you have $80.00 and your sister has $145.00. You decide to save $6.00 of your allowance each week, while your sister decide to spend her whole allowance plus $7.00 each week. How long will it be before you have as much money as your sister.
Answer by checkley77(12844)

80+6w=145-7w
6w+7w=145-80
13w=65
w=65/13
w=5 weeks they will have the same amount of money.
Proof:
80+6*5=145-7*5
80+30=145-35
110=110